A manufacturer of Christmas decorations has in 20x8 produced 1,000,000 large Christmas balls and 2,000,000 small
Christmas balls. Both products use the same raw materials and the production process
requires only minor adjustments when changing product. The company uses equivalence calculation for
to calculate the cost of the products.
The Christmas decoration manufacturer has calculated that a small Christmas ball requires 40% of the raw materials that go with a
large Christmas ball, 60% of direct salary and 50% of the indirect manufacturing costs involved in
production of large Christmas balls. The supplement for sales and administration is calculated as a percentage supplement
on manufacturing cost.
Total costs in 20x8 were:

Direct materials NOK 450,000
Direct salary NOK 660,000
Indirect manufacturing costs NOK 900,000
Sales and administration costs NOK 600,000

What is the cost of a small Christmas ball?
Answer #1
Cost of Small Christmas Ball
Particulars Cost Percentage
Direct Material 450000.00 40% Let cost of large balls= x, cost of small balls=0.4x
x+0.4x=450000, x= 321429
Cost of small balls= 128571
Salary 660000 60% Let cost of large balls= x, cost of small balls=0.6x
x+0.5x=660000, x= 412500
Cost of small balls= 247500
Indirect mfr cost 900000 50% Let cost of large balls= x, cost of small balls=0.5x
x+0.5x=900000, x= 600000
Cost of small balls= 300000
Total manufacturing cost of small balls= 128571+247500+300000= $676071
Total mfr cost of large balls= 321429+412500+600000= $1333929
Percentage for sales&admin cost= 676071/1333929= 50.68%
Sales & admin cost 600000 50.68% Let cost of large balls= x, cost of small balls=0.5068x
x+0.5068x=600000, x= 398195
Cost of small balls= 201805
Total cost of small christmas balls= 676071+201805= $877876
Unit cost= 877876/2000000= $0.44